# Robust Distributed Estimation: Extending Gossip Algorithms to Ranking and Trimmed Means

### v6

@@ C.1 Proof of Theorem 1
Let us assume that $\mathcal{G}=(V,E)$ is connected and non bipartite. Then, for $\mathbf{R}(t)$ defined in GoRank, we have that for all $k\in[n]$ :
@@
$|\mathbb{E}[R_{k}(t)]-r_{k}|\leq\frac{1}{ct}\cdot\sigma_{n}\left(r_{k}\right),$ (12)
where $c=\frac{\lambda_{2}}{|E|}$, with $\lambda_{2}$ being the second smallest eigenvalue of the graph Laplacian (spectral gap), and $\sigma_{n}(r_{k})=n\tilde{\sigma}_{n}\left(\frac{r_{k}-1}{n}\right)$, with $\tilde{\sigma}_{n}(x)=\sqrt{nx(1-x)}$.
@@
To prove Theorem 1, we will primarily use points (e) and (f) from Lemma.
@@
Denote $\tilde{\mathbf{W}}_{1}\triangleq\mathbf{W}_{1}-\frac{\mathbf{1}_{n}\mathbf{1}_{n}^{\top}}{n}$. Noticing that $\mathbf{h}_{k}^{\top}\frac{\mathbf{1}_{n}\mathbf{1}_{n}^{\top}}{n}\mathbf{e}_{k}=\bar{h}_{k}$ and that $\tilde{\mathbf{W}}_{1}^{s}=\mathbf{W}_{1}^{s}-\frac{\mathbf{1}_{n}\mathbf{1}_{n}^{\top}}{n}$ for $10\leq s\leq t-$ 1, we have
@@ C.2 Proof of Theorem 2
Let us assume that $\mathcal{G}=(V,E)$ is connected and non bipartite. Then, we have that for all $k\in[n]$, for any $t>0$ :
$\mathbb{E}[|R_{k}(t)-r_{k}|]\leq\sqrt{\frac{3}{ct}}\cdot\sigma_{n}(r_{k})\enspace,$ (18)
where $c$ and $\sigma_{n}(\cdot)$ are defined in Theorem 1.
@@
Thus, to prove Equation 18, a convergence result for the variance term $\mathbb{E}\left[\left(R_{k}(t)-r_{k}\right)^{2}\right]$ or equivalently $\mathbb{E}\left[\left(R^{\prime}_{k}(t)-\bar{h}_{k}\right)^{2}\right]$ is required. This result is formalized in the following lemma.
@@
$\mathbb{E}\left[\left(R_{k}(t)-r_{k}\right)^{2}\right]\leq\frac{3}{ct}\cdot\sigma_{n}\left(r_{k}\right)^{2},$
@@
where for any $11\leq s\leq t-$ 1, $\mathbf{W}(s:)=\mathbf{W}(s)\ldots\mathbf{W}(1)$ and $\mathbf{W}(0:)=\mathbf{I}_{n}$.
For any $t\geq 1$, let us define $\tilde{\mathbf{W}}_{1}(t)$ as $\tilde{\mathbf{W}}_{1}(t)\triangleq\mathbf{W}_{1}(t)-\frac{\mathbf{1}_{n\times n}}{n}$. Note that $\mathbf{W}_{1}(t)$ are a real symmetric matrices and that $\mathbf{1}_{n}$ is always an eigenvector of such matrices. Therefore, for any $10\leq s\leq t-$ 1, one has: $\tilde{\mathbf{W}}_{1}(s:)=\mathbf{W}_{1}(s:)-\frac{\mathbf{1}_{n\times n}}{n}$. Using this decomposition in (19) yields
@@ D.1 Proof of Lemma 1
Let $\mathbf{R}(t)$ and $\mathbf{W}(t)$ be defined as in Algorithm 1 and Algorithm 2, respectively. For all $k\in[n]$ and $t>0$, we have:
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\frac{3}{ct}\cdot\frac{\sigma_{n}(r_{k})^{2}}{\gamma_{k}^{2}(1-2\alpha)},$ (20)
where $\gamma_{k}=\min(\left|r_{k}-a\right|,\left|r_{k}-b\right|)\geq\frac{1}{2}$ with $a=\lfloor\alpha n\rfloor+\frac{1}{2}$ and $b=n-\lfloor\alpha n\rfloor+\frac{1}{2}$ being the endpoints of interval $I_{n,\alpha}$. The constants $c$ and $\sigma_{n}(\cdot)$ are as defined in Theorem 2.
@@
$\left|\mathbb{P}(R_{k}(t)\in I_{n,\alpha})-\mathbb{I}_{\{r_{k}\in I_{n,\alpha}\}}\right|\leq\mathbb{P}(|R_{k}(t)-r_{k}|\geq\gamma_{k})\leq\frac{3}{ct}\cdot\frac{\sigma_{n}\left(r_{k}\right)^{2}}{\gamma_{k}^{2}},$ (21)
@@ D.2 Proof of Theorem 3
where $\mathbf{K}=\left[\frac{\sigma_{n}(r_{k})^{2}}{\gamma_{k}^{2}}\right]_{k=1}^{n}$ and $c$, $\sigma_{n}(r_{k})$ and $\gamma_{k}$ are constants defined in Lemma 1.
@@
where $C_{i}={\frac{3}{c}}\cdot\frac{\sigma_{n}(r_{i})^{2}}{\gamma_{i}^{2}(1-2\alpha)}$ comes from Lemma 1 and we denote $\mathbf{C}=[C_{1},\ldots,C_{n}]$. To bound $\mathbf{R}(t)$, we decompose it using an intermediate time step $T=t-{\log(t)/c_{2}}>0$, where $c_{2}=1-\lambda_{2}(2)>0$. Let $T^{*}=\min\left\{t>1\,\middle|\,c_{2}t>\log(t)\right\}$. We obtain for all $t>T^{*}$,
@@
Moreover, we have $\|\mathbf{C}\odot\mathbf{X}\|={\frac{{3}}{c(1-2\alpha)}}\|\mathbf{K}\odot\mathbf{X}\|$ where $\mathbf{K}=\left[\frac{\sigma_{n}(r_{k})^{2}}{\gamma_{k}^{2}}\right]_{k=1}^{n},$ which completes the proof. ∎
@@ Appendix E Robustness Analysis: Breakdown Point of the Partial Trimmed Mean
where $T=\frac{4}{c}\log\left(\frac{n}{\varepsilon\delta}\right)$ denote a propagation time with $c$ being the connectivity constant and $\varepsilon:=\tau/\max_{i}|X_{i}|$ denote a tolerance parameter. Moreover, we define $K(\delta):=\frac{\sqrt{3}\sigma_{n}(m)}{\left(1-\frac{1}{n}\right)\sqrt{c}\delta}$.
@@
$\tilde{\varepsilon}^{*}(t)\geq\frac{\max\left(\left\lfloor m+\frac{1}{2}-\frac{c_{n}(m)}{\delta_{1}\sqrt{ct}}\right\rfloor,0\right)}{n},$
where we define $m=\lfloor n\alpha\rfloor$ and $c_{n}(m)=\sqrt{3}\sigma_{n}(m)$. As $t\to\infty$, note that $p=\lfloor\frac{1}{2}+m\rfloor=m$, which allows us to recover the breakdown point of the $\alpha$-trimmed mean.
@@
$p_{k}(t)\leq\frac{\sqrt{3}\sigma_{n}(r_{k})}{(a-r_{k})\sqrt{ct}}\leq\frac{\sqrt{3}\sigma_{n}(m)}{(a-r_{k})\sqrt{ct}},$
@@
$p_{k}(t)\leq\frac{c_{n}(m)}{(a-r_{k})\sqrt{ct}}\leq\delta.$
where $c_{n}(m)=\sqrt{3}\sigma_{n}(m)$ and obtain
$p=\max\left(\left\lfloor\frac{1}{2}+m-\frac{c_{n}(m)}{\delta\sqrt{ct}}\right\rfloor,0\right).$
@@
Let $\delta_{2}>0$ be a confidence parameter and $\varepsilon>0$ a tolerance parameter. We define $\mathbf{Z}(\cdot)$ as the evolution of the standard mean estimate, initialized as $\mathbf{Z}(0)=\mathbf{X}$. Additionally, we introduce a perturbed estimate, $\mathbf{\tilde{Z}}(\cdot)$, which starts at $\mathbf{\tilde{Z}}(0)=\mathbf{X}+B\mathbf{e}_{j}$. At iteration $t$, we detect a mistake of magnitude $B$, we correct it by injecting $-B$, leading to the update $\mathbf{\tilde{Z}}(t)\leftarrow\mathbf{\tilde{Z}}(t)-B\mathbf{e}_{j}$. Then, it follows that, with probability at least $1-\delta_{2}$, for all $s\geq t+T$, for any node $i$,
@@
Proof. At iteration $t$, the perturbed estimate is $\mathbf{\tilde{Z}}(t)=\mathbf{W}_{2}(t{:})\mathbf{\tilde{Z}}(0)=\mathbf{Z}(t)+\mathbf{W}_{2}(t{:})B\mathbf{e}_{j}$ since $\mathbf{Z}(t)=\mathbf{W}_{2}(t:)\mathbf{Z}(0)$. Then, at iteration $s>t$, we inject $-B$ at node $j$ which gives the following: $\mathbf{\tilde{Z}}(s)=\mathbf{W}_{2}(s{:}t)\left[\mathbf{Z}(t)+\mathbf{W}_{2}(t{:})B\mathbf{e}_{j}-B\mathbf{e}_{j}\right]$. Thus, for any $s>t$,
@@
Following the proof in [3], by repeatedly conditioning, we obtain
@@ Appendix F Convergence Analysis of ClippedGossip
where $\boldsymbol{L}_{ij}:=(\boldsymbol{e}_{i}-\boldsymbol{e}_{j})(\boldsymbol{e}_{i}-\boldsymbol{e}_{j})^{\top}$ is the elementary Laplacian associated with edge $(i,j)$ and $\alpha_{ij}^{t}:=\frac{1}{2}\min\left(1,\tau/\|{x}_{i}^{t}-{x}_{j}^{t}\|\right)$ a constant that depends on the previous estimates. Moreover, we have $\boldsymbol{W}(t)\boldsymbol{W}(t)^{\top}=\boldsymbol{I}_{n}-2\alpha_{ij}^{t}(1-\alpha_{ij}^{t})\boldsymbol{L}_{ij}$. In the special case $\alpha_{ij}=1/2$, the matrix reduces to the standard averaging and is a projection matrix: $\boldsymbol{W}(t)\boldsymbol{W}(t)^{\top}=\boldsymbol{W}(t)$
@@
by the triangle inequality. Similarly, we derive for $k\neq i,j$, $\|x_{j}^{1}-x_{k}^{0}\|\leq m.$ Finally, since $10<1-2\alpha_{ij}<$ 1, we have
@@ G.1 Experiments Compute Resources
The experiments are run on a single CPU with 32 GB of memory. The execution time for each experiment is given in Table 1.
@@ G.3 Additional Experiments
Figures 4, 5, 6, and 7 extend the experiments presented in sections 3 and 4.
@@
Figure 4: Role of $\sigma_{n}(\cdot)$ for three different communication graphs
@@
(a) Complete Graph
(b) Watts-Strogatz Graph
(c) 2D Grid Graph
Figure 6: Role of $c_{n}(r_{k})/\gamma_{k}$ for three different communication graphs
@@ Appendix H Asynchronous Variant of GoRank
In this section, we propose an asynchronous variant of GoRank that operates without access to a global clock or a shared iteration counter. Instead, each node $k$ maintains a local counter $C_{k}$ which tracks the number of times it has participated in an update. Asynchronous GoRank, outlined in Algorithm 6 proceeds as follows. When node $k$ is selected, it increments $C_{k}$ and updates its local rank estimate using a running average, where the global iteration $t$ is replaced by $C_{k}$. As in the synchronous version of GoRank, selected nodes also exchange their auxiliary observations. Note that this asynchronous version is more efficient, as it significantly reduces the number of updates—performing only two updates per iteration instead of $n$.
Figure 8 shows a comparison of Asynchronous GoRank with our two other ranking algorithms: Synchronous GoRank and Baseline++. The results suggest that Asynchronous GoRank converges slightly faster than Synchronous GoRank, thereby outperforming it in both speed and efficiency.
@@
Appendix I Further Discussion and Future Work

### v7

@@ C.1 Proof of Theorem 1
Let $\mathbf{R}(t)$ be defined in GoRank. We have that for all $k\in[n]$ :
@@
$|\mathbb{E}[R_{k}(t)]-r_{k}|\leq\frac{1}{ct}\cdot\sigma_{k},$ (11)
where $c=\frac{\lambda_{2}}{|E|}$, with $\lambda_{2}$ being the second smallest eigenvalue of the graph Laplacian (spectral gap), and $\sigma_{k}=n\tilde{\sigma}_{n}\left(\frac{r_{k}-1}{n}\right)$, with $\tilde{\sigma}_{n}(x)=\sqrt{nx(1-x)}$.
@@
To prove Theorem 1, we will primarily use points (e) and (f) from ˜Lemma.
@@
Denote $\tilde{\mathbf{W}}_{1}\triangleq\mathbf{W}_{1}-\frac{\mathbf{1}_{n}\mathbf{1}_{n}^{\top}}{n}$. Noticing that $\mathbf{h}_{k}^{\top}\frac{\mathbf{1}_{n}\mathbf{1}_{n}^{\top}}{n}\mathbf{e}_{k}=\bar{h}_{k}$ and that $\tilde{\mathbf{W}}_{1}^{s}=\mathbf{W}_{1}^{s}-\frac{\mathbf{1}_{n}\mathbf{1}_{n}^{\top}}{n}$ for $0\leq s\leq t-1$, we have
@@ C.2 Proof of Theorem 2
Let all $k\in[n]$, and let $c$ and $\sigma_{k}$ be as defined in Theorem˜1. Then, for all $t\geq 1$, we have: $\mathbb{E}[\left|R_{k}(t)-r_{k}\right|^{2}]\leq\left(3/ct\right)\cdot\sigma_{k}^{2}\,.$ Consequently,
$\mathbb{E}\left[\left|R_{k}(t)-r_{k}\right|\right]\leq\sqrt{\frac{3}{ct}}\cdot\sigma_{k}\enspace.$
@@
Thus, to prove Theorem˜2, a convergence result for the variance term $\mathbb{E}\left[\left(R_{k}(t)-r_{k}\right)^{2}\right]$ or equivalently $\mathbb{E}\left[\left(R^{\prime}_{k}(t)-\bar{h}_{k}\right)^{2}\right]$ is required. This result is formalized in the following lemma.
@@
$\mathbb{E}\left[\left(R_{k}(t)-r_{k}\right)^{2}\right]\leq\frac{3}{ct}\cdot\sigma_{k}^{2},$
@@
where for any $1\leq s\leq t-1$, $\mathbf{W}(s:)=\mathbf{W}(s)\ldots\mathbf{W}(1)$ and $\mathbf{W}(0:)=\mathbf{I}_{n}$.
For any $t\geq 1$, let us define $\tilde{\mathbf{W}}_{1}(t)$ as $\tilde{\mathbf{W}}_{1}(t)\triangleq\mathbf{W}_{1}(t)-\frac{\mathbf{1}_{n\times n}}{n}$. Note that $\mathbf{W}_{1}(t)$ are a real symmetric matrices and that $\mathbf{1}_{n}$ is always an eigenvector of such matrices. Therefore, for any $0\leq s\leq t-1$, one has: $\tilde{\mathbf{W}}_{1}(s:)=\mathbf{W}_{1}(s:)-\frac{\mathbf{1}_{n\times n}}{n}$. Using this decomposition in (17) yields
@@ D.1 Proof of Lemma 1
Let $\mathbf{R}(t)$ and $\mathbf{W}(t)$ be defined as in Algorithm˜1 and Algorithm˜2, respectively. For all $k\in[n]$ and $t>0$, we have:
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\frac{3}{ct}\cdot\frac{\sigma_{k}^{2}}{\gamma_{k}^{2}(1-2\alpha)},$ (18)
where $\gamma_{k}=\min(\left|r_{k}-a\right|,\left|r_{k}-b\right|)\geq\frac{1}{2}$ with $a=\lfloor\alpha n\rfloor+\frac{1}{2}$ and $b=n-\lfloor\alpha n\rfloor+\frac{1}{2}$ being the endpoints of interval $I_{n,\alpha}$. The constants $c$ and $\sigma_{n}(\cdot)$ are as defined in Theorem˜2.
@@
$\left|\mathbb{P}(R_{k}(t)\in I_{n,\alpha})-\mathbb{I}_{\{r_{k}\in I_{n,\alpha}\}}\right|\leq\mathbb{P}(|R_{k}(t)-r_{k}|\geq\gamma_{k})\leq\frac{3}{ct}\cdot\frac{\sigma_{k}^{2}}{\gamma_{k}^{2}},$ (19)
@@ D.2 Proof of Theorem 3
where $\mathbf{K}=\left[\frac{\sigma_{k}^{2}}{\gamma_{k}^{2}}\right]_{k=1}^{n}$ and $c$, $\sigma_{k}$ and $\gamma_{k}$ are constants defined in Lemma 1.
@@
where $C_{i}={\frac{3}{c}}\cdot\frac{\sigma_{k}^{2}}{\gamma_{i}^{2}(1-2\alpha)}$ comes from Lemma 1 and we denote $\mathbf{C}=[C_{1},\ldots,C_{n}]$. To bound $\mathbf{R}(t)$, we decompose it using an intermediate time step $T=t-{\log(t)/c_{2}}>0$, where $c_{2}=1-\lambda_{2}(2)>0$. Let $T^{*}=\min\left\{t>1\,\middle|\,c_{2}t>\log(t)\right\}$. We obtain for all $t>T^{*}$,
@@
Moreover, we have $\|\mathbf{C}\odot\mathbf{X}\|={\frac{{3}}{c(1-2\alpha)}}\|\mathbf{K}\odot\mathbf{X}\|$ where $\mathbf{K}=\left[\frac{\sigma_{k}^{2}}{\gamma_{k}^{2}}\right]_{k=1}^{n},$ which completes the proof. ∎
@@ Appendix E Robustness Analysis: Breakdown Point of the Partial Trimmed Mean
where $T=\frac{4}{c}\log\left(\frac{n}{\varepsilon\delta}\right)$ denote a propagation time with $c$ being the connectivity constant and $\varepsilon:=\tau/\max_{i}|X_{i}|$ denote a tolerance parameter. Moreover, we define $K(\delta):=\frac{c_{m}}{\left(1-\frac{1}{n}\right)\sqrt{c}\delta}$ where $c_{m}=\sqrt{3}n^{3/2}\cdot\phi((m-1)/n)$ with $\phi:u\in(0,1)\to\sqrt{u(1-u)}$.
@@
$\tilde{\varepsilon}^{*}(t)\geq\frac{\max\left(\left\lfloor m+\frac{1}{2}-\frac{c_{m}}{\delta_{1}\sqrt{ct}}\right\rfloor,0\right)}{n},$
where we define $m=\lfloor n\alpha\rfloor$ and $c_{m}=\sqrt{3}n^{3/2}\cdot\phi((m-1)/n)$. As $t\to\infty$, note that $p=\lfloor\frac{1}{2}+m\rfloor=m$, which allows us to recover the breakdown point of the $\alpha$-trimmed mean.
@@
$p_{k}(t)\leq\frac{\sqrt{3}\sigma_{n}(r_{k})}{(a-r_{k})\sqrt{ct}}\leq\frac{c_{m}}{(a-r_{k})\sqrt{ct}},$
@@
$p_{k}(t)\leq\frac{c_{m}}{(a-r_{k})\sqrt{ct}}\leq\delta,$
and obtain
$p=\max\left(\left\lfloor\frac{1}{2}+m-\frac{c_{m}}{\delta\sqrt{ct}}\right\rfloor,0\right).$
@@
[…]
@@
Proof. At iteration $t$, the perturbed estimate is $\mathbf{\tilde{Z}}(t)=\mathbf{W}_{2}(t{:})\mathbf{\tilde{Z}}(0)=\mathbf{Z}(t)+\mathbf{W}_{2}(t{:})B\mathbf{e}_{j}$ since $\mathbf{Z}(t)=\mathbf{W}_{2}(t:)\mathbf{Z}(0)$. Then, at iteration $s>t$, we inject B-B at node $j$ which gives the following: $\mathbf{\tilde{Z}}(s)=\mathbf{W}_{2}(s{:}t)\left[\mathbf{Z}(t)+\mathbf{W}_{2}(t{:})B\mathbf{e}_{j}-B\mathbf{e}_{j}\right]$. Thus, for any $s>t$,
@@
Following the proof in Boyd et al. [2006], by repeatedly conditioning, we obtain
@@ Appendix F Convergence Analysis of ClippedGossip
where $\boldsymbol{L}_{ij}:=(\boldsymbol{e}_{i}-\boldsymbol{e}_{j})(\boldsymbol{e}_{i}-\boldsymbol{e}_{j})^{\top}$ is the elementary Laplacian associated with edge $(i,j)$ and $\alpha_{ij}^{t}:=\frac{1}{2}\min\left(1,\tau/\|{x}_{i}^{t}-{x}_{j}^{t}\|\right)$ a constant that depends on the previous estimates. Moreover, we have $\boldsymbol{W}(t)\boldsymbol{W}(t)^{\top}=\boldsymbol{I}_{n}-2\alpha_{ij}^{t}(1-\alpha_{ij}^{t})\boldsymbol{L}_{ij}$. In the special case $\alpha_{ij}=1/2$, the matrix reduces to the standard averaging and is a projection matrix: $\boldsymbol{W}(t)\boldsymbol{W}(t)^{\top}=\boldsymbol{W}(t)$ .
@@
by the triangle inequality. Similarly, we derive for $k\neq i,j$, $\|x_{j}^{1}-x_{k}^{0}\|\leq m.$ Finally, since $0<1-2\alpha_{ij}<1$, we have
@@ G.1 Experiments Compute Resources
The experiments are run on a single CPU with 32 GB of memory.
The execution time for each experiment is less than 30 minutes (except for the large-scale experiments). The details of a few experiments are given in Table˜1.
@@ G.3 Additional Experiments
Figures 4, 5, and 6 extend the experiments presented in sections 3 and 4.
@@
Figure 4: Role of $\phi$ for three different communication graphs
@@ Appendix H Asynchronous Variant of GoRank
In this section, we propose an asynchronous variant of GoRank that operates without access to a global clock or a shared iteration counter. Instead, each node $k$ maintains a local counter $C_{k}$ which tracks the number of times it has participated in an update. Asynchronous GoRank, outlined in Algorithm˜6 proceeds as follows. When node $k$ is selected, it increments $C_{k}$ and updates its local rank estimate using a running average, where the global iteration $t$ is replaced by $C_{k}$. As in the synchronous version of GoRank, selected nodes also exchange their auxiliary observations. Note that this asynchronous version is more efficient, as it significantly reduces the number of updates—performing only two updates per iteration instead of $n$.
Figure˜7 shows a comparison of Asynchronous GoRank with our two other ranking algorithms: Synchronous GoRank and Baseline++. The results suggest that Asynchronous GoRank converges slightly faster than Synchronous GoRank, thereby outperforming it in both speed and efficiency.
@@
Note that the convergence analysis in the asynchronous case is more complex because it requires analyzing the ratio of two statistics. Nonetheless, though technically more demanding, it is possible to derive convergence rate bounds using Taylor expansions techniques. We will carry out such an analysis in an extension to this work.
Appendix I Large-scale Experiments
To demonstrate the scalability of our method on large networks, we repeated the experiments from Fig. (c) in Sections 3.2 and 4.3, originally conducted with $n=500$, on larger networks with $n=1000$ and $n=5000$.
For the experiments related to Section 3.2, the results lead to the same conclusions: Figure˜8 shows that GoRank continues to achieve a low error (<0.1<0.1), and the overall performance trends of the other ranking algorithms remain similar on the Watts–Strogatz graph but are significantly worse on the 2D grid graph. These additional results further reinforce GoRank’s scalability and robustness, particularly in comparison to the other methods.
For the experiments related to Section 5.4, Figure˜9 shows that GoTrim + GoRank continues to converge efficiently to the trimmed mean, even on larger 2D grid graphs, clearly outperforming the corrupted mean. In contrast, GoTrim + Baseline++ converges significantly more slowly on the 2D grid (and does not even outperform the corrupted mean), although it remains efficient on Watts–Strogatz graphs.
(a) Watts-Strogatz with $n=1000$
(b) 2D grid with $n=1000$
(c) 2D grid with $n=5000$
Figure 8: Comparison of the performance of gossip algorithms for ranking across large networks.
(a) Watts-Strogatz with $n=1000$
(b) 2D grid with $n=1000$
(c) 2D grid with $n=5000$
Figure 9: Comparison of the performance for trimmed means estimation across large networks.
Appendix J Robustness to Network Disruptions
While robustness to data contamination is important, the robustness of our proposed algorithms to network disruptions (e.g., edge/node failures, network partitioning) is equally crucial in real-world applications.
One natural extension is to introduce a fixed probability of failure for each node or edge (see [2]). This would modify the edge sampling: instead of a uniform distribution, the algorithm would sample edge $e$ with probability $p_{e}$. In our analysis, this change corresponds to replacing the normalized Laplacian $L/|E|$ with a weighted sum $\sum_{e}p_{e}L_{e}$, where $\sum_{e}p_{e}<1$ due to edge failures and $L_{e}$ correspond to the elementary Laplacians. The spectral gap, which governs the convergence rate, would now correspond to the connectivity constant $\tilde{c}$ of this weighted Laplacian instead of $c=\lambda_{2}/|E|$. For instance, if each edge fails independently with probability 0.5, then $p_{e}=1/2|E|$ for all $e$, and the effective spectral gap becomes $\tilde{c}=\lambda_{2}/(2|E|)<c$, indicating a slower convergence rate. We emphasize that as long as $p_{e}>0$ for all $e$, the weighted graph remains connected and non-bipartite if the original graph was.
Another interesting scenario is network partitioning. One way to model this is by designing graphs composed of tightly connected clusters with only a few inter-cluster edges that are prone to failure. In such a setup, the connectivity constant would degrade significantly, and we expect the convergence rate to reflect this bottleneck. We generated a graph consisting of 500 nodes organized into three well-connected clusters, with only five inter-cluster edges. As expected, we observed a very low connectivity constant $c=1.86\times 10^{-6}$ and the convergence behavior of this graph is similar to that of a 2D grid graph, which is consistent with the low overall connectivity.
Appendix K Experiments on Sparse Graphs
An interesting question is whether our algorithms performance well on sparse graphs. First, we note that the Watts-Strogatz and 2D grid graphs used in our main experiments are already quite sparse, with both containing fewer than 1000 edges. However, we would like to emphasize that what primarily governs convergence behavior is not sparsity per se, but graph connectivity. For example, a dense graph composed of loosely connected clusters may have poor connectivity, while a sparse graph like a 3-regular graph can exhibit strong connectivity properties. A cycle graph, in addition to being sparse, has very low connectivity and serves as a useful pathological case. To better illustrate the relationship between topology and performance, we conducted additional experiments using three different sparse graphs: 1) Watts-Strogatz graph (connectivity $c=3.31\times 10^{-4}$, 1̃000 edges), 2) Cycle graph ($c=3.16\times 10^{-7}$, 499 edges), 3) 3-regular graph ($c=2.24\times 10^{-4}$, 750 edges). All experiments were run for $15\times 10^{4}$ iterations. For GoRank, both the Watts-Strogatz and 3-regular graphs achieved fast convergence (absolute error below 0.05), consistent with their strong connectivity properties typical of expander graphs. The cycle graph, as expected, performed worse due to its very low connectivity, but still achieved an error below 0.1. For GoTrim, using the same corruption model as in Fig. 5b, the Watts-Strogatz and 3-regular graphs again performed very well, with errors close to 0. However, the cycle graph exhibited much slower convergence: after $8\times 10^{4}$ iterations, the absolute error remained above 40. Doubling the number of iterations reduced the error to below 40. While this is indeed slow, it is important to note that performance remains better than the corrupted mean, and such a topology is highly atypical in real-world sensor networks. In practice, such scenarios would likely require gossip algorithms specifically designed for cycle graphs. These experiments support the conclusion that GoTrim still performs well on sparse graph with good connectivity properties.
(a) GoRank on sparse graphs
(b) GoTrim on expander graph
(c) GoTrim on cycle graph
Figure 10: Comparison of the performance of GoRank and GoTrim on different sparse graphs.
Appendix L Further Discussion and Future Work